OpenStudy (anonymous):

Needs help understanding how to solve this question! It really confuses me! sqrt(c)-sqrt(d)/sqrt(c)+sqrt(d)

OpenStudy (campbell_st):

one problem is that is isn't an equation as the statement isn't equal to anything

OpenStudy (anonymous):

Ok the instructions say to rationalize the denominator. Assume all expressions under radicals represent positive numbers.

OpenStudy (anonymous):

When you have a sum of square roots on the bottom, you can multiply the bottom by the conjugate... Multiply the top and bottom by sqrt(c) - sqrt(d) to get the answer.

OpenStudy (anonymous):

so I am multiplying and one will cancel out and since I am multiply by sqrt(c)-sqrt(d) the top cancels?

OpenStudy (campbell_st):

rationalising the denominator means multiplying by something that will give a rational number.... you always multiply by 1 which is written in another form in your question use the denominator with the opposite sign \[1 = (\sqrt{c} - \sqrt{d})/(\sqrt{c} -\sqrt{d})\] so the problem is now \[(\sqrt{c} - \sqrt{d})/(\sqrt{c} + \sqrt{d}) \times (\sqrt{c} - \sqrt{d})/(\sqrt{c} - \sqrt{d})\] this simplifies to \[(\sqrt{c} - \sqrt{d})^2 /(c - d)\] the radicals are removed from the denominator

OpenStudy (anonymous):

Ok...I see now...thanks so much!

OpenStudy (anonymous):

One more question...when typing in my answer, are the ( ) needed?